Find the number of two-digit positive integers whose digits total 7. I know the answer is 7 (16, 25, 34, 43, 52, 61, 70) but my teacher wants me to show it in an algebraic formula. How do I do that? This question is not from a textbook.
Answer by checkley77(12569)

X+Y=7
MAKE UP A TABLE WITH X=0,1,2,3,4,5,6 & 7 AND SOLVE FOR Y.
0+Y=7
Y=7 OR 07
1+Y=7
Y=7-1
Y=6 OR 16
2=Y=7
Y=7-2
Y=5 OR 25
3=Y=7
Y=7-3
Y=4 OR 34
5+Y=7
Y=7-5
Y=2 OR 52
6+Y=7
Y=7-6
Y=1 OR 61
7+Y=7
Y=7-7
Y=0 OR 70